### v2

# Fast Bilateral Teleoperation and Imitation Learning Using Sensorless Force Control via Accurate Dynamics Model

Abstract
In recent years, the advancement of imitation learning has led to increased interest in teleoperating low-cost manipulators to collect demonstration data. However, most existing systems rely on unilateral control, which only transmits target position values. While this approach is easy to implement and suitable for slow, non-contact tasks, it struggles with fast or contact-rich operations due to the absence of force feedback. This work demonstrates that fast teleoperation with force feedback is feasible even with force-sensorless, low-cost manipulators by leveraging 4-channel bilateral control. Based on accurately identified manipulator dynamics, our method integrates nonlinear terms compensation, velocity and external force estimation, and variable gain corresponding to inertial variation. Furthermore, using data collected by 4-channel bilateral control, we show that incorporating force information into both the input and output of learned policies improves performance in imitation learning. These results highlight the practical effectiveness of our system for high-fidelity teleoperation and data collection on affordable hardware.
Figure 1: Fast bilateral teleoperation on force sensorless low cost manipulator
Keywords: Teleoperation, Force Control, Imitation Learning
1 Introduction
In recent years, influenced by the remarkable achievements of Deep Learning in fields such as images and language, research into the application of Deep Learning to robot motion generation has attracted attention. In particular, imitation learning  [1] , which uses human motion as the correct answer in supervised learning, is attracting attention because it enables high-sample-efficiency learning in robot learning, which involves significant costs for data collection on actual robots. In particular, neural network models that have been trained using imitation learning with extremely large amounts of human motion data  [2] [3] [4] are called “robot foundation models” in analogy to pre-training models for natural language processing, attracting a lot of attention.
In this context, there is an increasing number of examples of low-cost hardware being used to collect teaching data for robot learning, like ALOHA series [5][6][7] and GELLO [8] . This is thought to be due to the fact that a large number of robots are needed to collect a large amount of data, and it is more important to keep the cost per robot down, and also because the policies learned sometimes result in undesirable behavior, and low-cost, relatively small hardware can reduce the risk of danger or failure.
However, most of these systems used unilateral control, which only transmits target position values from the leader robot to the follower robot [9]. This method is easy to implement and provides good operability in low-speed, low-contact operations, but it does not provide force feedback and struggles with contact-rich tasks. In fact, it also has difficulties in high-speed operations. On the other hand, there has been an increase in cases where bilateral control, which enables two-way information exchange between the leader robot and the follower robot and provides force feedback, has been adopted. There are various types of bilateral control. In some cases, force reflecting type bilateral control is used, in which external forces from followers are generated feedforward-style on the leader side, like FACTR [10] . However, this method is known to have stability issues, and in many cases, stability is ensured by adding a damping term [10][11], but this tends to impair operability. On the other hand, 4-channel bilateral control is a method in which both the leader and follower transmit their respective positions and forces to each other and perform position control and force control in parallel. Theoretically, a 4-channel architecture can achieve the ideal response of synchronization of position and force perfectly [12][13]. However, to perform position control and force control with opposite objectives simultaneously in a single robot, making them non-interfering requires accurate dynamics models and observation of position, velocity, and external forces. Because the dynamics of manipulators with six or more degrees of freedom involve many parameters and extremely complex functions, obtaining accurate parameters and performing real-time calculations has long been impractical. As a practical implementation method, an accelerated control system was proposed that combines a simplified dynamics model based on linear models of each joint with disturbance estimation (including external force and model error) using a disturbance observer [14] and compensation using the estimated disturbance. This enables high performance even when a certain amount of model error exists [15][16][17][18][19]. This acceleration control-based 4-channel bilateral control system has also been studied for application to imitation learning. Sakaino et al. trained a Long-Short-Term Memory (LSTM) [20] using data collected with four-channel bilateral control and conducted experiments to perform contact-rich tasks [21][22][23]. Kobayashi et al. published Alpha-$\alpha$ and Bi-ACT series [24][25][26][27] that are a combination of a low-cost 4-channel bilateral control system Alpha-$\alpha$ and the method to train Action Chunking with Transformer (ACT) [5] using data collected by 4-channel bilateral control.
However, when using low-cost hardware, factors such as slow control cycles, insufficient resolution of rotary encoders, and significant backlash prevent rapid compensation for model errors from functioning effectively, thereby limiting effectiveness. In addition, the problem of modeling errors caused by using each axis linear approximation model has been largely solved by using hardware with high reduction ratios. However, this solution may not work as well when using direct drive motors or pseudo-direct drive motors, which have been increasingly used in recent years. On the other hand, with the increase in computing resources and the evolution of algorithms, parameter identification and real-time calculations of the dynamics of multi-degree-of-freedom manipulators are becoming practical.
In this study, we focused on 4-channel bilateral control combining a parameter-identified, accurate nonlinear manipulator dynamics model with a disturbance observer. By effectively utilizing feedforward based on accurate dynamics models, high tracking performance and operability can be achieved even with low-cost hardware that cannot use high feedback gains. We compared the proposed method with other remote control techniques, such as unilateral control, position-symmetric control, and force-feedback control, as well as with conventional 4-channel bilateral control with simplified dynamics models, in high-speed motions, and confirmed the improvement in control performance achieved by the proposed method. As a result, we have demonstrated the ability to teach high-difficulty tasks involving multiple contacts, such as quickly turning a nut or peeling a cucumber, using a low-cost manipulator. Additionally, we conducted imitation learning using the collected data and verified the task success rate.
Figure 2: Block diagram of controller and observer
2 Methodology
2.1 4-channel Bilateral Control
We will explain a 4-channel bilateral control system that uses a rigid serial link model. The block diagram is shown in Fig. 4. Although many studies on 4-channel bilateral control have been published in the past, the major feature of the control system used in this study is that it utilizes parameter identification of the nonlinear dynamics of the manipulator’s rigid body model and incorporates compensation for the nonlinear term, as well as state estimation based on this.
In general, assuming that each link is a rigid body, the dynamics of the manipulator can be expressed as a rigid serial link model.
$\bm{M}(\bm{\theta})\ddot{\bm{\theta}}+\bm{C}(\bm{\theta},\dot{\bm{\theta}})\dot{\bm{\theta}}+\bm{D}\dot{\bm{\theta}}+\bm{g}(\bm{\theta})=\bm{\tau}_{ref}+\bm{\tau}_{ext}$ (1)
where $\bm{M}(\bm{\theta})$ is the inertial matrix, $\bm{C}(\bm{\theta},\dot{\bm{\theta}})\dot{\bm{\theta}}$ is the centrifugal and Coriolis forces, $\bm{D}$ is the viscous friction coefficient, $\bm{g}(\bm{\theta})$ is the gravitational forces and $\bm{\tau}_{ref},\bm{\tau}_{ext}$ are the reference and external torques of each joint, respectively. After compensating, the dynamics can be written as follows:
$\bm{\tau}_{u}$ $\coloneqq\bm{\tau}_{ref}-\tilde{\bm{C}}(\bm{\theta},\hat{\dot{\bm{\theta}}})\hat{\dot{\bm{\theta}}}-\tilde{\bm{D}}\hat{\dot{\bm{\theta}}}-\tilde{\bm{g}}(\bm{\theta})$ (2)
$\bm{\tau}_{u}$ $=\bm{M}(\bm{\theta})\ddot{\bm{\theta}}-\bm{\tau}_{ext}$ (3)
where $\bm{\tau}_{u}$ is the reference torque other than compensation, $\tilde{\bigcirc}$ denotes parameter-identified manipulator dynamics functions, and $\hat{\bigcirc}$ denotes values estimated by an observer, which will be explained later. Conventionally, it was difficult to perform bilateral control using strict dynamic models such as (2) and (3). We propose to conduct parameter identification using the regressor matrix based on OpenSYMORO [28], which is software to derive efficient robot dynamics parameters, and the least squares method. Here, we consider the coordinate transformation of the joint angular velocity of the leader and follower into differences and averages.
$\begin{bmatrix}\dot{\bm{\theta}}_{-}\\ \dot{\bm{\theta}}_{+}\\ \end{bmatrix}=\begin{bmatrix}\bm{I}&-\bm{I}\\ \frac{1}{2}\bm{I}&\frac{1}{2}\bm{I}\\ \end{bmatrix}\begin{bmatrix}\dot{\bm{\theta}}_{l}\\ \dot{\bm{\theta}}_{f}\\ \end{bmatrix}\coloneqq\bm{J}_{+-}\begin{bmatrix}\dot{\bm{\theta}}_{l}\\ \dot{\bm{\theta}}_{f}\\ \end{bmatrix}$ (4)
where $\dot{\bm{\theta}}_{-}$ and $\dot{\bm{\theta}}_{+}$ are difference and average of the joint angler velocity. This transformation allows us to think of position and force control independently. In other words, if the same amount of velocity is applied to both the leader and the follower in opposite directions, only the difference in the joint angle will change, not the average. Conversely, if the same amount of velocity is applied to both the leader and the follower in the same direction, only the average joint angle will change, not the difference. Ideally, these do not influence each other, i.e., they are orthogonal, so control systems can be designed separately for each. Since the coordinate transformation matrix is time-invariant, the coordinate transformation of the angle and angular acceleration can be described by the same matrix $\bm{J}_{+-}$. The dynamics of the difference and the average of the joint angles are as follows.
$\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{u}$ $=\begin{bmatrix}\bm{M}_{l}(\bm{\theta}_{l})&\bm{0}\\ \bm{0}&\bm{M}_{f}(\bm{\theta}_{f})\\ \end{bmatrix}\begin{bmatrix}\ddot{\bm{\theta}}_{l}\\ \ddot{\bm{\theta}}_{f}\\ \end{bmatrix}-\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}=\begin{bmatrix}\bm{M}_{l}(\bm{\theta}_{l})&\bm{0}\\ \bm{0}&\bm{M}_{f}(\bm{\theta}_{f})\\ \end{bmatrix}\bm{J}_{+-}^{-1}\begin{bmatrix}\ddot{\bm{\theta}}_{-}\\ \ddot{\bm{\theta}}_{+}\\ \end{bmatrix}-\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}$ (5)
For teleoperation, the desired dynamics for the joint angle difference are fixed, which means heavy weight, significant friction, and high stiffness. On the other hand, the desired dynamics for the average joint angle are light and have no spring or damper. The desired dynamics can be written as a spring-mass-damper system like impedance control [29] on transformed coordinates [16][30] as below:
$\begin{bmatrix}\bm{\tau}_{-}\\ \bm{\tau}_{+}\\ \end{bmatrix}_{ext}=\bm{J}_{+-}^{-\top}\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}=\begin{bmatrix}\frac{1}{2}\bm{I}&-\frac{1}{2}\bm{I}\\ \bm{I}&\bm{I}\\ \end{bmatrix}\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}=\begin{bmatrix}\bm{M}_{-}&\bm{0}\\ \bm{0}&\bm{M}_{+}\\ \end{bmatrix}_{des}\begin{bmatrix}\ddot{\bm{\theta}}_{-}\\ \ddot{\bm{\theta}}_{+}\\ \end{bmatrix}+\begin{bmatrix}\bm{D}_{-}&\bm{0}\\ \bm{0}&\bm{0}\\ \end{bmatrix}_{des}\begin{bmatrix}\dot{\bm{\theta}}_{-}\\ \dot{\bm{\theta}}_{+}\\ \end{bmatrix}+\begin{bmatrix}\bm{K}_{-}&\bm{0}\\ \bm{0}&\bm{0}\\ \end{bmatrix}_{des}\begin{bmatrix}\bm{\theta}_{-}\\ \bm{\theta}_{+}\\ \end{bmatrix}$ (6)
where $\bigcirc_{des}$ are desired dynamics parameters, and $\bm{\tau}_{-,ext},\bm{\tau}_{+,ext}$ are external forces corresponding to the difference and average of the joint angle. Since the system has position control and force control on orthogonal axes, the system can also be interpreted as position/force hybrid control [31] on transformed coordinates [17]. The control law can be derived by substituting the desired dynamics into the transformed dynamics:
$\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{u}$ $=\begin{bmatrix}\bm{M}_{l}(\bm{\theta}_{l})&\bm{0}\\ \bm{0}&\bm{M}_{f}(\bm{\theta}_{f})\\ \end{bmatrix}\begin{bmatrix}\frac{1}{2}\bm{I}&\bm{I}\\ -\frac{1}{2}\bm{I}&\bm{I}\\ \end{bmatrix}\begin{bmatrix}-\bm{M}_{-,des}^{-1}\bm{D}_{-,des}\dot{\bm{\theta}}_{-}-\bm{M}_{-,des}^{-1}\bm{K}_{-,des}\bm{\theta}_{-}\\ \bm{M}_{+,des}^{-1}\bm{\tau}_{+}\\ \end{bmatrix}-\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}$ (7)
$=\begin{bmatrix}\bm{M}_{l}(\bm{\theta}_{l})&\bm{0}\\ \bm{0}&\bm{M}_{f}(\bm{\theta}_{f})\\ \end{bmatrix}\begin{bmatrix}\bm{K}_{p}(\bm{\theta}_{f}-\bm{\theta}_{l})+\bm{K}_{d}(\dot{\bm{\theta}}_{f}-\dot{\bm{\theta}}_{l})+\bm{K}_{f}(\bm{\tau}_{l,ext}+\bm{\tau}_{f,ext})\\ \bm{K}_{p}(\bm{\theta}_{l}-\bm{\theta}_{f})+\bm{K}_{d}(\dot{\bm{\theta}}_{l}-\dot{\bm{\theta}}_{f})+\bm{K}_{f}(\bm{\tau}_{l,ext}+\bm{\tau}_{f,ext})\\ \end{bmatrix}-\begin{bmatrix}\bm{\tau}_{l}\\ \bm{\tau}_{f}\\ \end{bmatrix}_{ext}$ (8)
where $\bm{K}_{p}=\frac{1}{2}\bm{M}_{-,des}^{-1}\bm{K}_{-,des}$, $\bm{K}_{d}=\frac{1}{2}\bm{M}_{-,des}^{-1}\bm{D}_{-,des}$, and $\bm{K}_{f}=\bm{M}_{+,des}^{-1}$. Finally, the control law for each manipulator can be as follows:
$\bm{\tau}_{ref}$ $=\tilde{\bm{M}}(\bm{\theta})\left\{\bm{K}_{p}(\bm{\theta}_{des}-\bm{\theta})+\bm{K}_{d}(\dot{\bm{\theta}}_{des}-\hat{\dot{\bm{\theta}}})+\bm{K}_{f}(\bm{\tau}_{des}+\hat{\bm{\tau}}_{ext})\right\}-\hat{\bm{\tau}}_{ext}+\tilde{\bm{C}}(\bm{\theta},\hat{\dot{\bm{\theta}}})\hat{\dot{\bm{\theta}}}+\tilde{\bm{D}}\hat{\dot{\bm{\theta}}}+\tilde{\bm{g}}(\bm{\theta})$ (9)
where $\bm{\theta}_{des},\dot{\bm{\theta}}_{des},\bm{\tau}_{des}$ are opposite side states between leader and follower. The result can be understood as a combination of the computed torque method, position PD control, and force control.
2.2 Velocity and External Force Observer
To implement the above 4-channel bilateral control, the joint angles, joint angular velocities, and external forces of the leader and follower are required. However, the manipulator used in this study is equipped only with 12-bit rotary encoders to measure joint angles. Therefore, the joint angular velocities and external forces must be estimated by an observer. Although many methods for estimating the velocity of a manipulator have been proposed [32][33], here we introduce an observer that can estimate both the velocity and the external forces simultaneously.
Assuming that the differential of the external forces remain zero ($\frac{d}{dt}\left\{\bm{M}(\bm{\theta})^{-1}\bm{\tau}_{ext}\right\}=\bm{0}$), the dynamics of the manipulator can be described in the following state equation form.
$\frac{d}{dt}\begin{bmatrix}\bm{\theta}\\ \dot{\bm{\theta}}\\ \bm{M}(\bm{\theta})^{-1}\bm{\tau}_{ext}\\ \end{bmatrix}=\begin{bmatrix}\bm{0}&\bm{I}&\bm{0}\\ \bm{0}&\bm{0}&\bm{I}\\ \bm{0}&\bm{0}&\bm{0}\\ \end{bmatrix}\begin{bmatrix}\bm{\theta}\\ \dot{\bm{\theta}}\\ \bm{M}(\bm{\theta})^{-1}\bm{\tau}_{ext}\\ \end{bmatrix}+\begin{bmatrix}\bm{0}\\ \bm{I}\\ \bm{0}\\ \end{bmatrix}\bm{M}(\bm{\theta})^{-1}\bm{\tau}_{u},\quad\bm{y}$ $=\begin{bmatrix}\bm{I}&\bm{0}&\bm{0}\\ \end{bmatrix}\begin{bmatrix}\bm{\theta}\\ \dot{\bm{\theta}}\\ \bm{M}(\bm{\theta})^{-1}\bm{\tau}_{ext}\\ \end{bmatrix}$ (10)
Minimal-order observers are often used to minimize the phase delay of external force estimation in disturbance observers [14][34][35]. Although several methods have been proposed for constructing minimal-order observers, here we explain a specific, simplified method. First, assuming that velocity can be observed, we will construct a full-order observer using a state-space model with only velocity and external force.
$\frac{d}{dt}\begin{bmatrix}\hat{\dot{\bm{\theta}}}\\ \tilde{\bm{M}}(\bm{\theta})^{-1}\hat{\bm{\tau}}_{ext}\\ \end{bmatrix}=\begin{bmatrix}\bm{0}&\bm{I}\\ \bm{0}&\bm{0}\\ \end{bmatrix}\begin{bmatrix}\hat{\dot{\bm{\theta}}}\\ \tilde{\bm{M}}(\bm{\theta})^{-1}\hat{\bm{\tau}}_{ext}\\ \end{bmatrix}+\begin{bmatrix}\bm{I}\\ \bm{0}\\ \end{bmatrix}\tilde{\bm{M}}(\bm{\theta})^{-1}\bm{\tau}_{u}+\begin{bmatrix}2\zeta\omega_{c}\bm{I}\\ \omega_{c}^{2}\bm{I}\\ \end{bmatrix}(\frac{d}{dt}\bm{\theta}-\hat{\dot{\bm{\theta}}})$ (11)
Here, the observer gain is set using $\omega_{c}$, corresponding to the cut-off angular frequency in the frequency domain. The critical point is that inertia is not a constant but a function of the joint angle vector. Conventional disturbance observers use constants for inertia, and the effects of modeling errors are included in the estimated external force. Therefore, it was assumed that inertia could be nominalized by feedback compensation of the external force. However, the extent to which modeling errors can be compensated depends on how high the cut-off frequency (observer gain) can be set. In practice, when using control equipment capable of achieving high-precision rotary encoders and high sampling frequencies, it is possible to set a fairly high cut-off frequency, and by combining fast external force estimation and compensation for modeling errors, it is practically feasible to ignore modeling errors. However, in the case of low-cost hardware like the one used here, a high cut-off frequency cannot be set, so the error in inertia has a significant impact that cannot be ignored. By pre-identifying the nonlinear dynamics of the manipulator with 38 parameters, including the inertial matrix, and using this to always employ accurate inertial parameters, it is possible to improve control performance while maintaining a low cost.
Furthermore, using the identified inertia justifies the use of a velocity observer. In many studies of 4-channel bilateral control, pseudo-differentiation ($\frac{\omega_{c}s}{s+\omega_{c}}$) is used for velocity estimation, but pseudo-differentiation is known to cause estimation delays [32]. Nevertheless, pseudo-differentiation is widely used because the inertia is not accurately known, and methods that can calculate using only angle response values without using inertia are preferred. If accurate inertia can be obtained, it becomes possible to use not only the information from the angular response values but also predictions from the torque reference values, thereby improving the delay in speed estimation. Improving the delay in speed estimation allows the D gain of the position PD control to be set higher, and setting the D gain higher enables the P gain to be set higher as well, leading to a significant improvement in position control performance.
Converting this to a transfer function expression using the Laplace transform, we can write it as follows.
$\hat{\dot{\bm{\theta}}}$ $=\frac{s}{s+2\zeta\omega_{c}}\frac{1}{s}\ddot{\bm{\theta}}_{ref}+\frac{2\zeta\omega_{c}}{s+2\zeta\omega_{c}}s\bm{\theta}\qquad\left(\ddot{\bm{\theta}}_{ref}\coloneqq\bm{M}(\bm{\theta})^{-1}(\bm{\tau}_{u}+\hat{\bm{\tau}}_{ext})\right)$ (12)
$\hat{\bm{\tau}}_{ext}$ $=\tilde{\bm{M}}(\bm{\theta})\frac{\omega_{c}^{2}}{s^{2}+2\zeta\omega_{c}s+\omega_{c}^{2}}(s^{2}\bm{\theta}-\tilde{\bm{M}}(\bm{\theta})^{-1}\bm{\tau}_{u})$ (13)
We implemented the algorithm using a bilinear transformation ($s=\frac{2}{T}\cdot\frac{1-z^{-1}}{1+z^{-1}}$).
3 Experiments
((a)) vs. other teleoperation methods
((b)) ablation study
Figure 3: Results of teleoperation comparison
3.1 Teleoperation Performance
We implemented the proposed 4-channel bilateral control on CRANE-X7, a 7 DoF low-cost manipulator manufactured by RT Corporation. We compared teleoperation performance by repeatedly swinging the joint closest to the base as quickly as possible at a 90-degree angle 10 times. The results are shown in Figure 3. We compared three other teleoperation techniques and conducted three ablation studies.
vs. Unilateral Control. Unilateral control is a teleoperation method that controls the position of the follower by using the leader’s position as the target position. The biggest drawback of unilateral control is the lack of force feedback for the operator, though position tracking performance is also an issue. Compared to the 4-channel type, which uses position and force control for both the leader and follower robots, unilateral control uses only position control for the follower, resulting in limited tracking performance at the same gain. As a result, the position error became larger than that of the proposed method.
vs. Symmetric Position Type Bilateral Control. Symmetric position type bilateral control is a method in which both the leader and follower control their positions based on each other’s positions as targets. It is known that this method causes the operation to feel very heavy due to the force that acts to keep the current position unchanged. In this case, since the manipulator was relatively lightweight, the heavy operation problem did not seem so serious. On the other hand, there was a difference in the position tracking error. In the case of a 4-channel type, the large operating force applied to the leader when the leader changes direction is also applied to the follower by force control, enabling rapid directional changes. On the other hand, in a symmetric position type that relies solely on position control, large feedback was not generated unless a large position error occurred, making it difficult to follow rapid directional changes.
vs. Force Feedback Type Bilateral Control. Force feedback type bilateral control is a method in which the follower side performs position error feedback and the leader side performs force error feedback. Like unilateral control, this type also relies solely on the position control of the follower to reduce position error. Consequently, its position tracking performance was lower than that of the 4-channel type and symmetric position type at the same gain. However, due to the force control on the leader side, its tracking performance is slightly higher than that of unilateral control.
vs. 4ch Bilateral Control using fixed inertia. We compared the results obtained using fixed inertial parameters to those obtained when inertial variations were considered. Although control systems with fixed inertia (fixed gain) are simple and easy to stabilize, and they are often used in manipulator control, their performance depends on posture and is not constant. The results of the experiment showed that considering inertial variations significantly reduced position errors.
vs. 4ch Bilateral Control w/o centrifugal & coriolis compensation. Centrifugal and Coriolis forces are often ignored because their effects are negligible in cases of low speed, high friction, and high reduction ratios. In this experiment, however, the forces had a greater impact than traditional industrial manipulators due to lower friction (due to hardware characteristics and friction compensation). When the joint closest to the base rotated quickly, the centrifugal force caused the other joints to move quickly without the operator touching them. This is shown in the figure, where the force applied by the operator was slightly larger.
vs. 4ch Bilateral Control using pseudo differential. Velocity estimation using the pseudo-differential method shows vibration tendencies due to estimation delays, and cannot use high gains. In the figure, observing the motion after stopping confirms that high-frequency vibrations occured when using velocity estimation based on the pseudo-differential method. To use it in practice, it will be necessary to significantly decrease the gain of position control.
Table 1: Success rate of dual arm pick-and-place tasks
Action Chunking with Transformer
[…]
pick place pick place pick place
10mm 0/5 0/5 5/5 0/5 5/5 5/5
20mm 0/5 0/5 5/5 5/5 5/5 5/5
30mm 5/5 0/5 5/5 5/5 5/5 5/5
40mm 5/5 0/5 5/5 4/5 5/5 5/5
50mm 5/5 0/5 5/5 0/5 5/5 5/5
Total 15/25 0/25 25/25 14/25 25/25 25/25
Table 2: Success rate of imitation learning tasks
ACT
[…]
Nut Turning 0/5 5/5 5/5
Cucumber Peeling 0/5 3/5 2/5
3.2 Imitation Learning
We experimented to collect demonstration data using the proposed 4-channel bilateral teleoperation method and perform imitation learning. We used Action Chunking with Transformer (ACT) [5] for neural network architecture. The combination of ACT and 4-channel bilateral control is called Bi-ACT [24], and has already been validated on Pick-and-Place and Put-in-drawer tasks. With 4-channel bilateral control, data on the position and force of the leader and follower can be obtained. In addition to using all of this data, we also performed learning using only position information and force information only on the input side. We investigated the effects of using force information on both the input and output sides.
It is worth noting that a control system utilizing position and force command values was employed during data collection. Since the control system differs between data collection and autonomous operation, a decrease in success rate was expected when force information is not used in the policy output. However, differences in control systems or dynamics between data collection and autonomous operation are common when using data collection methods other than teleoperation, and this is not uncommon in imitation learning. Additionally, since the dataset’s quality can significantly affect imitation learning results, it is preferable to compare using the same dataset. Considering these points, we conducted this comparison.
Task1: Dual-arm Pick-and-Place. We performed a task of grasping blocks of different widths between two arms. We prepared five types of blocks, ranging from 10 mm to 50 mm in width, and collected a total of 10 demonstration data sets, two for each type. The blocks are shown in Fig. 8, and the results are shown in Table 2. As a result, when no force information was used for either input or output, picking was not possible in cases with small widths of 10 mm and 20 mm. In cases with relatively large widths of 30 mm to 50 mm, picking was possible once, but problems such as dropping the object immediately or the angle changing while holding it occurred, and the object could not be placed in the correct position. When force information was used in the input, the success rate of picking improved significantly, with all trials resulting in successful picking. However, instability after picking persisted, and placement failures occurred frequently in cases where the width was farthest from the average of 10mm and 50mm. When both input and output used force information, both pick and place were successful in all trials. The stability of holding after picking improved, with no instances of dropping midway or misalignment.
Task2: Nut Turning. We tried the nut turning task as a task that requires more force and speed. The results are shown in the first row of Table 2. As a result, when no force was applied to either the input or output, the experiment was completely unsuccessful, but when force was applied to the input, the experiment was successful regardless of whether force was applied to the output. When no force was applied to the input, the hand often stopped at the position where it overlapped with the nut on the bird’s-eye view image, and it seemed difficult to touch the nut due to insufficient force information.
Task3: Cucumber Peeling. As a task involving handling irregularly shaped objects, we tried peeling cucumbers with a peeler. The results are shown in the third row of Table 2. As a result, when no force information was used for either input or output, the experiment was completely unsuccessful. When force information was used for input, the experiment was successful in 3/5 cases, and when force information was used for both input and output, the experiment was successful in 2/5 cases. The main failure case was when the system transitioned to the finished state before touching the cucumber or peeler. This is thought to be because there was not much difference between the image at the start of the task and the image after the task was completed.
4 Conclusion
In this study, we demonstrated that a low-cost manipulator can achieve high-speed teleoperation with force feedback without force sensors by incorporating a high-precision manipulator dynamics model into a four-channel bilateral teleoperation system. We compared the proposed method to three teleoperation methods, such as unilateral control, symmetric position type, and force feedback type, with a 4-channel system that does not utilize manipulator dynamics. We confirmed that the proposed method demonstrates higher position and force tracking performance during high-speed operations. Furthermore, we collected demonstration data using the proposed teleoperation method. To investigate the effectiveness of utilizing force information in both the input and output of the policy, we conducted imitation learning. As a result, adding force input showed a significant improvement in success rate regardless of the task. On the other hand, the effect of force output depended on the task. Future challenges include improving data collection and optimization methods for parameter identification to more modern methods [36][37], theoretical examination of variable gain control systems, and the investigation of more advanced neural network architectures for imitation learning other than force control systems, as well as compatibility with a wider range of tasks other than those attempted in this study.
5 Limitation
This study has several limitations. Although the theory supports cases where the leader and follower have different inertia matrices, verification has only been performed using the same inertia matrix. Although stability can be compensated for under the assumption of valid feedback linearisation, this has not been proven in cases where nonlinearity persists due to modelling errors. Furthermore, the data used for parameter identification was collected using unilateral control, which is not theoretically substantiated. Regarding imitation learning, application to neural network architectures other than ACT has not been performed, and the quantity of data and number of trials is limited.
Acknowledgments
This work was supported by JSPS KAKENHI Grant Number 24K00905, 24KJ0503, JST PRESTO Grant Number JPMJPR24T3, and JST ALCA-Next Japan, Grant Number JPMJAN24F1. This study was based on the results obtained from the JPNP20004 project subsidized by the New Energy and Industrial Technology Development Organization (NEDO).